OpenStudy (anonymous):

what is the derivative of 1/x ?

OpenStudy (saifoo.khan):

\[\Large \frac{1}{x} \to x^{-1}\]

OpenStudy (anonymous):

is it not x^-2 ?

OpenStudy (hba):

@saifoo.khan She just asked x^4 * 0 =0 whether it was true or false lol :D

OpenStudy (anonymous):

lol she did

OpenStudy (saifoo.khan):

Where? I can't read the question here.

OpenStudy (hba):

@Veneria Maybe if your question was 1/x^2 then it would go up and become x^-2

OpenStudy (saifoo.khan):

@Veneria : Now find it's derivative.

OpenStudy (hba):

d/dx x^a = ax^a-1

OpenStudy (anonymous):

you gotta do first principles, there's no other way

OpenStudy (hba):

Please derivate it ?

OpenStudy (anonymous):

how?

OpenStudy (hba):

\[d/dx x^a = ax^{a-1} \]